OpenStudy (anonymous):

PLEASE HELP!!! Solve 4^x = 12 a.) x ≈ 0.27 b.) x ≈ 0.56 c.) x ≈ 1.79 d.) x ≈ 19.93

OpenStudy (mathstudent55):

None of the answers are correct.

OpenStudy (anonymous):

i copied and pasted the question and answers directly form the site.

OpenStudy (anonymous):

sry but these are all not it

OpenStudy (mathstudent55):

That's better.

OpenStudy (mathstudent55):

\(4^x = 12\) Take log of both sides: \(\log{4^x} = \log{12}\) Use law of logs: \(x \log {4} = \log{12} \) Solve for x

OpenStudy (mathstudent55):

One of the answers is correct.

OpenStudy (anonymous):

Ah yes! I must have missed the power sign! [problem with computers and maths!] Yes, one of the answers is correct!

OpenStudy (mathstudent55):

@sarahusher You didn't miss the power sign. It was added later.

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\] and a calculator

OpenStudy (anonymous):

looks to be about 1.79...